Tom and Jan together earn $84,326 per year. If Tom earns 12,468 more per year than Jan, then how much does each of them earn per year?
Answer by AnlytcPhil(1806)


Tom and Jan together earn $84,326 per year. If Tom earns 12,468 more per year
than Jan, then how much does each of them earn per year?
`
>>...Tom and Jan together earn $84,326 per year...<<
`
Translation: T + J = 84326
`
If Tom earns 12,468 more per year than Jan, then how much does each of them
earn per year?<<>>
`
Translation: T = J + 12468
`
So you have to solve this system of equations.
`
T + J = 84326
T = J + 12468
`
Can you solve this system by substitution? If not post again.
`
Answer: T = $48397, J = $35929
`
Edwin